OpenStudy (anonymous):

What is the solution to the equation? –4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12

OpenStudy (shaadykhamlichi):

www.wolframalpha.com

OpenStudy (anonymous):

q=12

OpenStudy (anonymous):

and thats your answer

OpenStudy (shaadykhamlichi):

why r u telling answers ???

OpenStudy (anonymous):

Simplifying -4q + 8 = -2q + -16 Reorder the terms: 8 + -4q = -2q + -16 Reorder the terms: 8 + -4q = -16 + -2q Solving 8 + -4q = -16 + -2q Solving for variable 'q'. Move all terms containing q to the left, all other terms to the right. Add '2q' to each side of the equation. 8 + -4q + 2q = -16 + -2q + 2q Combine like terms: -4q + 2q = -2q 8 + -2q = -16 + -2q + 2q Combine like terms: -2q + 2q = 0 8 + -2q = -16 + 0 8 + -2q = -16 Add '-8' to each side of the equation. 8 + -8 + -2q = -16 + -8 Combine like terms: 8 + -8 = 0 0 + -2q = -16 + -8 -2q = -16 + -8 Combine like terms: -16 + -8 = -24 -2q = -24 Divide each side by '-2'. q = 12 Simplifying q = 12

OpenStudy (anonymous):

you gave him a website...

OpenStudy (shaadykhamlichi):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

like thats any better

OpenStudy (shaadykhamlichi):

website is not the answer

OpenStudy (anonymous):

ok thank you for showing ur work

OpenStudy (anonymous):

np

OpenStudy (anonymous):

nice monkey pic

OpenStudy (anonymous):

plz fan me and medal

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np